OpenStudy (anonymous):

Can someone check/verify my answer? Write the expression as either the sine, cosine, or tangent of a single angle. sin(pi/2)cos(pi/7)+cos(pi/2)sin(pi/7) My work: cos(A-B) = cos A cos B + sin A sin B A = pi/2 B = pi/7 cos(pi/2-pi/7)=2pi/14

OpenStudy (psymon):

Looks like you picked the wrong identity. sinAcosB + sinBcosA is the sum of sines formula.

OpenStudy (anonymous):

ohhhh so I should use the sin one then.

OpenStudy (psymon):

Yep yep. So same angles, but sin and addition.

OpenStudy (anonymous):

Okay so: sinAcosB+sinBcosA=sin(A+B) A=pi/2 B=pi/7 sin(pi/2)cos(pi/7)+cos(pi/2)sin(pi/7) = sin 9pi/14

OpenStudy (psymon):

Yep, looks good to me.

OpenStudy (anonymous):

Yay! Thanks!

OpenStudy (psymon):

Sure ^_^